OpenStudy (anonymous):

The temperature of an object is- 40 degree celsius. what will be its temperature on the fahrenheit scale?

OpenStudy (anonymous):

http://www.manuelsweb.com/temp.htm here's a good link.

OpenStudy (anonymous):

i nned to understand the concept so it would be nice of you to explain.

OpenStudy (anonymous):

oh, okay. since the guy defined the freezing point of water to be 32F, and the boiling point of water to be 212F, (212-32)F=(100-0)C 180F=100C 5F=9C so, since there should be equal increase of volume of the thermometer liquid(the rise in mercury/alcohol level is the same), 5 degrees increase of Fahrenheit scale will equal 9 degrees increase of Celcius 5F=9C 1C=5/9 F -40C=-40(5/9) F=22.22F

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Fahrenheit the pic in here should be helpful

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/File:Thermometer_CF.svg

OpenStudy (anonymous):

thankyou, i appreciate it

OpenStudy (anonymous):

you're welcome :)